OpenStudy (anonymous):

is 40y-4y²+10-y the correct expansion for (4y+1)(10-y)?

OpenStudy (anonymous):

(4y+1)(10+y) = 40y + 4y^2+10-y

OpenStudy (chaise):

Nope - do you need help with FOIL? (4y+1)(10-y) = (4y*10)-(4y*y)+(1*10)+(1*-y) = 40y - 4y^2 +10 -y = 39y - 4y^2 +10 = -4y^2+39y+1-

OpenStudy (anonymous):

Well it was still right but I could have gone one more step down right?

OpenStudy (chaise):

Oh, yes, you were right, it just wasn't the simplified version. Whenever you see terms that can be simplified you should always simplify them. For example in an equation you wouldn't leave it as 8+2, you would write 10. Same goes for variables. 40y - y = 39y.

OpenStudy (anonymous):

right right. Thanks alot bros.